Table 1. Smoking status before onset of the present illness, lung cancer cases and matched controls with other diseases, Great Britain, 1948-1952.

Cases

Controls

Cigarette smoker

1,350

1,296

Non-Smoker

7

61

Total

1,357

1,357

Question 11 (2 points) Calculate the odds of smoking among the cases.
Answer #1

Cigarette smoker cases = 1,350

Total Cases= 1,357

Odds ratio = probability/ (1-probability) = proportion /(1-proportion)

(1,350/1,357) / (7/1,357) = 1,350/7 = 192.9:1

According to the given case group patients would smoke 193 times probably.